50 ml of a 0.1m koh solution is added to 25ml of 0.2m hcl. the volume of the resulting solution was made upto 500 ml. now 0.745 gram of solid kcl is added to solution . what is the concentration of cl- in the final solution . what volume of 0.1 m agno3 would be requried for complete precipitaion of cl - ions as silver chloride
1
Expert's answer
2019-12-03T08:25:55-0500

Solution.

"KOH + HCl = KCl + H2O"

"n(KOH) = C(KOH) \times V(KOH)"

"n(HCl) = C(HCl) \times V(HCl)"

n(KOH) = 0.005 mol

n(HCl) = 0.005 mol

n(KOH) = n(HCl) = n(KCl) = 0.005 mol

"m(KCl) = 0.005 \times 74.5513 = 0.373 \ g"

"m'(KCl) = 0.373 + 0.745 = 1.118 \ g"

"KCl = K^+ + Cl^-"

C(KCl) = C(Cl-)

"C'(KCl) = \frac{m'(KCl)}{M(KCl) \times V}"

"C'(KCl) =\frac{1.118}{74.5513 \times 0.5} = 0.03 M"

C(Cl-) = 0.03 M

"AgNO3 + KCl = AgCl + KNO3"

"C(AgNO3) \times V(AgNO3) = C(KCl) \times V(KCl)"

"V(AgNO3) = \frac{C(KCl) \times V(KCl)}{C(AgNO3)}"

"V(AgNO3) = \frac{0.03 \times 0.5}{0.1} = 0.15 L = 150 \ ml"

Answer:

C(Cl-) = 0.03 M

V(AgNO3) = 150 ml = 0.15 L